Charles has to pay $133 per year for the next 16 years. If the appropriate discount rate is 0.135, what would it cost today to settle his debt?
Answer #1

Information provided:

Yearly payment= $133

Time= 16 years

Discount rate= 0.135*100= 13.50%

The question is solved by computing the present value.

Enter the below in a financial calculator to compute the present value:

PMT= 133

N= 16

I/Y= 13.50

Press the CPT key and PV to compute the present value.

The value obtained is 855.2927.

Therefore, it will cost Charles $855.29 to settle his debts today.

In case of any query, kindly comment on the solution